OpenStudy (anonymous):

Justify each step. 7a + (4 + 12a) = 12a + 4

OpenStudy (anonymous):

7a + 4 + 12a = 12a + 4 -> just open ( ) 19a + 4 = 12a + 4 -> adding all the 'a' on the left side now subtract 12a from both sides to get : 19a - 12a + 4 = 12a - 12a + 4 7a + 4 = 4 subtract 4 from both sides to get 7a + 4 -4 = 4 -4 7a = 0 obviously a = 0